# Trajectory Balance with Asynchrony: Decoupling Exploration and Learning for Fast, Scalable LLM Post-Training

### v1

@@ 5.5 Discovery of high-reward samples via scaling search
Beyond improvements at a given level of compute, we also observe improvements when we scale the amount of total compute, showing TBA’s promise for large-scale distributed RL. In our asynchronous setup, we find that adding more searchers consistently improves the attack success rate and diversity for RT (see Figure 5, right). This improvement likely stems from having more searchers exploring different regions of the solution space simultaneously, enabling more effective discovery of high-reward samples. Moreover, asynchronous updates introduce opportunities for beneficial randomness, and thus potential expansion of the search coverage in the combinatorial space of language. Interestingly, we also find some evidence for scaling’s helpfulness in PFT (see Figure 7).
@@ 6 Discussion
In this work, we introduced TBA, a novel post-training method for large language models (LLMs) that combined an off-policy reinforcement learning (RL) objective with distributed asynchronous search. By decoupling searcher and trainer nodes, our approach enabled efficient distributed training and avoided bottlenecks, leading to significant performance gains in post-training tasks such as mathematical reasoning, automated red-teaming, and RLHF. We expect that our highly parallelizable and performant framework for RL post-training can be extended to other valuable tasks, including self-improvement training (Zelikman et al., 2022; Hu et al., 2024; Gulcehre et al., 2023), search-based reasoning (Wan et al., 2024), as well as the emerging paradigm of training LLMs to reason with RL (Guo et al., 2025).
Towards a multi-agent search system
Our approach can potentially be extended to a multi-agent search system aimed at identifying multiple diverse modes within the large combinatorial space of language. Our current searchers do not explicitly target different regions, but we do show that multiple distributed searchers can enhance exploration and populate a global replay buffer with varied experiences. By evolving this concept into an explicit multi-agent system, we can encourage each agent to explore distinct regions in a divide-and-conquer manner. This cooperative strategy allows agents to effectively identify and report local modes to a centralized replay buffer. The off-policy post-training agent can then learn from the diverse samples stored in this centralized buffer, benefiting from a wider range of experiences. We consider this a promising future direction.
Improving local credit assignment
The trajectory balance objective can suffer from high gradient variance as it operates on the trajectory level. We addressed this by sampling more responses per query. Future work can leverage learning partial energy functions (Madan et al., 2023; Zhang et al., 2025) to balance bias and variance during policy updates.
Acknowledgements
The authors acknowledge funding from CIFAR, NSERC, IVADO, and Samsung.
@@ Acknowledgements
Appendix A Experimental Details
@@
For MR and PFT, we implement TBA by building on the RLOO trainer class of the Hugging Face TRL library (von Werra et al., 2020). For MR and PFT, we make modifications to the baseline hyperparameters as shown in Table 3 and Table 4, respectively.
Our RT implementation of TBA built on the TB trainer used in Lee et al. (2025), and we largely utilize their code’s default hyperparameters, with differences noted in Appendix A.3.
A.1 GSM8K Mathematical Reasoning (MR)
All baselines and TBA results use the same starting point, a RhoMath-1B (Lin et al., 2024) model SFTed on GSM8K training data by Kazemnejad et al. (2024) – “realtreetune/rho-1b-sft-GSM8K” on Hugging Face. The baseline model achieves $40.3$% test accuracy. For the VinePPO baseline, training time is estimated using their reported 380 seconds per training step and 650 steps for GSM8K training (Kazemnejad et al., 2024).
@@ A.1 GSM8K Mathematical Reasoning (MR)
Most-On-Policy Sampling Probability 0.95 $m$ in Section 5.4
@@
Table 3: TBA Training Hyperparameters for the GSM8K MR task.
In Figure 1, TBA uses the settings in Table 3 with a few modifications to shorten training time by an additional $30$%, down to 82 minutes on 4xA100 GPUs. In particular, we observed in initial testing (see Appendix B) that using the hyperparameters in Table 3 led to no improvement in performance for the final 300 steps. Thus, we shrank the training duration from 1000 steps to 700 (98000 episodes with batch size 140). Additionally, we made the following modifications to attempt to reduce the variance of the shortened run: 350 stable learning rate steps (down from 450), 0.014 Beta Initial Value (up from 0.012). We ran this experiment three times – obtaining performances of $55.8$%, $53.9$%, and $54.1$% – and reported the mean accuracy $54.6$% in Figure 1.
@@
The 700-step result we show in Figure 1 has standard error $0.6$%, which is a little more variance than what we observe in the original 1000 step setup (the blue line in the bottom left plot of Figure 6 shows the mean and standard errors for the 1000 step runs). Future work could further explore variance reduction approaches/hyperparameters for shorter runs (and for TBA/RL more generally).
One way to deal with variance is to choose a checkpoint based on a held out validation set, a strategy used to produce the VinePPO result (Kazemnejad et al., 2024). We do not use this approach but note that our results would likely benefit significantly ($\approx 1$%) from it. In particular, each of our runs tends to produce (at some point during training) a higher performance than the performance at the final step – this is expected if you consider that the model performance varies around the average value it converges to towards the end of training (e.g., see again the blue line in the bottom left plot of Figure 6). Despite its being lower than the maximum performance achieved at any point during training, we report this final step performance, which is what a user of TBA could expect to obtain without applying techniques like early-stopping.
@@
Most-On-Policy Sampling Probability 0.5 $m$ in Section 5.4
@@
A.2 TL;DR Preference Fine Tuning (PFT)
@@
All of our PFT results were run in 32-bit precision and without DeepSpeed, which was used by baselines we compared against (Noukhovitch et al., 2025). Lower precision and such training acceleration packages could further improve the speedups we show. Relatedly, for our $2.8$B runs, we used gradient checkpointing to fit more examples into a micro batch, which led to slowdowns at this scale (i.e., we only have a $3.8$x speedup over the baseline in this setting). We leave the optimization of our framework with appropriate packages to future work. Finally, we used a large number (10000) of initial completions in the buffer, and future work should confirm that a smaller number (e.g. 1000) works equally well – note that a small number worked well for GSM8K.
A.3 Automated Red Teaming (RT)
Unlike our MR and PFT implementations, our RT implementation uses the code of Lee et al. (2025) as a baseline and thus does not follow the Hugging Face TRL trainer style. We discuss hyperparameters in the context of their trajectory balance trainer approach below. We adopt their code largely as it is, with the exception that our TBA implementation uses larger replay buffers than Lee et al. (2025) to accommodate our scaled search for trajectories. Additionally, unlike Lee et al. (2025), we remove the oldest samples when the buffers fill up as opposed to the lowest reward samples.
For the Llama results in Table 5.4, our hyperparameter choices largely follow those of Lee et al. (2025). We train for 5000 steps with batch size 128. For temperature sampling, we use a low and high of 0.7 and 2.0, respectively. We use a reward schedule horizon of 1000. The language model schedule end is 1.2, and its schedule’s horizon is 2000. We prioritize sampling based on reward. We use Beta 0.05. We use sync period ($k$) 10. We use 6 searchers. We use most-on-policy probability ($m$) 0.5 and 0.6 in combination with maximum buffer sizes 150000 and 130000, respectively. By having a smaller maximum buffer size and larger $m$, the latter setting is expected to focus on more recently generated data and prioritize reward over diversity, which is what we observe in Table 5.4.
For the GPT2 results in Figure 5 and Table 5.4, our hyperparameter choices again largely follow those of Lee et al. (2025). We train for 5000 steps with batch size 128. We use Beta 0.05. We use sync period ($k$) 10. We use most-on-policy probability ($m$) 0.5. We cap the maximum buffer size at 100000 in all experiments; we additionally prevent the buffer from growing past its size as of step 4000 to encourage focusing on more recent data (larger most on policy probabilities $m$ may provide a similar effect). We test searcher counts 2, 6, 14, 30, 62.
@@ A.3 Automated Red Teaming (RT)
In Table 5.4, there is a slowdown as the number of searchers scales. We believe this is largely addressed by a newer version of our code that uses more efficient buffer communication, but we have not re-run these results yet to confirm this. In any case, developing more efficient TBA implementations is an interesting direction for future work given TBA’s ability to effectively leverage large-scale data generation.
Appendix B GSM8K Ablation Studies

### v2

@@ 5.5 Discovery of high-reward samples via scaling search
Figure 5: TBA reaches the RT diversity-toxicity Pareto frontier and improves as search is scaled. (Left) On the GPT-2 automated red-teaming task of lee2024learning, TBA produces results on the diversity vs. toxicity Pareto frontier in less training time. Baselines taken from lee2024learning. (Right) Each searcher uses one V100 GPU for generating attacks. We report means and standard errors from multiple runs of the automated red-teaming task with GPT-2 at each searcher/GPU count.
Beyond improvements at a given level of compute, we also observe improvements when we scale the amount of total compute, showing TBA’s promise for large-scale distributed RL. In our asynchronous setup, we find that adding more searchers consistently improves the attack success rate and diversity for RT (see Figure˜5, right).
This improvement likely stems from having more searchers exploring different regions of the solution space simultaneously, enabling more effective discovery of high-reward samples. Moreover, asynchronous updates introduce opportunities for beneficial randomness, and thus potential expansion of the search coverage in the combinatorial space of language. Interestingly, we also find some evidence for scaling’s helpfulness in PFT (see Figure 14).
Relatedly, in Table 6, we show we can trade attack toxicity for greater attack diversity by scaling the maximum buffer size to retain more off-policy data. This RT experiment uses Llama 3.2 1B models.
6 Scaling TBA for Larger Models
As described in Section 4, TBA allows us to test a simplified version of TBA and to train with larger models/contexts. Relative to TBA, TBA has fewer knobs to tune: it does not have a decay schedule for $\beta$, and PRIME-RL does not sample from a buffer, using a constant level of off-policyness instead. A TBA configuration entails a reference policy reset interval $\rho$, and a constant $\beta$. We use $\rho=50$, $\beta=0.005$. In Figure 6, we find further evidence that TBA can enhance async RL training of LLMs, particularly when data is highly off-policy. See Appendices C and D for more details and results.
Figure 6: TBA vs. Dr. GRPO. When training Qwen 2.5 7B base [qwen25] on the MATH dataset [hendrycksmath2021] or Qwen 3 4B Instruct on Countdown [stojanovski2025reasoninggymreasoningenvironments], we find that TBA performs well relative to state-of-the-art methods, especially in highly off-policy settings. See Appendix D for additional results and details.
@@ 7 Discussion
In this work, we introduced TBA, a novel post-training method for large language models (LLMs) that combined an off-policy reinforcement learning (RL) objective with distributed asynchronous search. By decoupling searcher and trainer nodes, our approach enabled efficient distributed training and avoided bottlenecks, leading to significant performance gains in post-training tasks such as mathematical reasoning, automated red-teaming, and RLHF. We expect that our highly parallelizable and performant framework for RL post-training can be extended to other valuable tasks, including self-improvement training [zelikman2022star, hu2024amortizing, gulcehre2023reinforced], search-based reasoning [wan2024alphazero], as well as the emerging paradigm of training LLMs to reason with RL [guo2025deepseek].
Limitations
The trajectory balance objective can suffer from high gradient variance as it operates on the trajectory level. We addressed this by sampling more responses per query. Future work can leverage learning partial energy functions [madan2023learninggflownetspartialepisodes, zhang2025lessons] to balance bias and variance during policy updates.
Broader Impact
Improving RL training strategies for LLMs can enhance their usefulness across domains but also carries risks associated with misuse, reward misspecification, and unintended generalization. Careful evaluation and responsible deployment are essential as these methods scale.
Acknowledgments and Disclosure of Funding
We thank Michael Noukhovitch and Haizhong Zheng for helpful feedback.
The authors acknowledge funding from CIFAR, NSERC, IVADO, and Samsung. MJ is supported by a FRQNT Doctoral Fellowship (https://doi.org/10.69777/366694).
@@
Appendix A TBA Gradient Analysis
Below, we reproduce for convenience the TBA loss $\mathcal{L}_{\text{TB}}^{\text{VarGrad}}$ from Eq. 5, then provide the gradient of the corresponding maximization objective $\mathcal{J}_{\text{TB}}$.
$\begin{split}\mathcal{L}_{\text{TB}}^{\text{VarGrad}}(\mathbf{B};\theta)=\frac{1}{BK}\sum_{i=1,j=1}^{i=B,j=K}\bigg(\text{STOP-GRAD}[\log\hat{Z}(\mathbf{x}^{(i)})]+\log\pi_{\theta}(\mathbf{y}^{(i,j)}\mid\mathbf{x}^{(i)})\\ -\log\pi_{\text{ref}}(\mathbf{y}^{(i,j)}\mid\mathbf{x}^{(i)})-\frac{1}{\beta}r_{\phi}(\mathbf{y}^{(i,j)};\mathbf{x}^{(i)})\bigg)^{2}.\end{split}$ (8)
We first multiply everything inside the square operation of Eq. 5 by 1-1, which is equivalent to multiplying the entire equation by $1$ and gives
$\begin{split}\mathcal{L}_{\text{TB}}^{\text{VarGrad}}(\mathbf{B};\theta)=\frac{1}{BK}\sum_{i=1,j=1}^{i=B,j=K}\bigg(\log\pi_{\text{ref}}(\mathbf{y}^{(i,j)}\mid\mathbf{x}^{(i)})-\log\pi_{\theta}(\mathbf{y}^{(i,j)}\mid\mathbf{x}^{(i)})+\\ \frac{1}{\beta}r_{\phi}(\mathbf{y}^{(i,j)};\mathbf{x}^{(i)})-\log\hat{Z}(\mathbf{x}^{(i)})\bigg)^{2}.\end{split}$ (9)
The gradient with respect to $\theta$ then becomes
$\begin{split}\nabla\mathcal{L}_{\text{TB}}^{\text{VarGrad}}(\mathbf{B};\theta)=\frac{1}{BK}\sum_{i=1,j=1}^{i=B,j=K}-2\bigg(\log\pi_{\text{ref}}(\mathbf{y}^{(i,j)}\mid\mathbf{x}^{(i)})-\log\pi_{\theta}(\mathbf{y}^{(i,j)}\mid\mathbf{x}^{(i)})+\\ \frac{1}{\beta}r_{\phi}(\mathbf{y}^{(i,j)};\mathbf{x}^{(i)})-\log\hat{Z}(\mathbf{x}^{(i)})\bigg)\nabla\log\pi_{\theta}(\mathbf{y}^{(i,j)}\mid\mathbf{x}^{(i)}),\end{split}$ (10)
and we can turn this into a maximization problem instead of a minimization problem by multiplying the right side by 1-1:
$\begin{split}\nabla\mathcal{J}_{\text{TB}}^{\text{VarGrad}}(\mathbf{B};\theta)=\frac{1}{BK}\sum_{i=1,j=1}^{i=B,j=K}2\bigg(\log\pi_{\text{ref}}(\mathbf{y}^{(i,j)}\mid\mathbf{x}^{(i)})-\log\pi_{\theta}(\mathbf{y}^{(i,j)}\mid\mathbf{x}^{(i)})+\\ \frac{1}{\beta}r_{\phi}(\mathbf{y}^{(i,j)};\mathbf{x}^{(i)})-\log\hat{Z}(\mathbf{x}^{(i)})\bigg)\nabla\log\pi_{\theta}(\mathbf{y}^{(i,j)}\mid\mathbf{x}^{(i)}).\end{split}$ (11)
We multiply the right hand side by $\frac{\beta}{2}$, rescaling the gradient without changing the minimizer. Also, we define $\bar{\text{KL}}^{(i)}=\frac{1}{K}\sum_{j=1}^{K}\hat{\text{KL}}^{(i,j)}=\frac{1}{K}\sum_{j=1}^{K}\log\pi_{\theta}(\mathbf{y}^{(i,j)}\mid\mathbf{x}^{(i)})-\log\pi_{\text{ref}}(\mathbf{y}^{(i,j)}\mid\mathbf{x}^{(i)})$ and define $\bar{r}^{(i)}=\frac{1}{K}\sum_{j=1}^{K}r^{(i,j)}=\frac{1}{K}\sum_{j=1}^{K}r_{\phi}(\mathbf{y}^{(i,j)};\mathbf{x}^{(i)})$ to rewrite Eq. 4 as $\log\hat{Z}(\mathbf{x}^{(i)})=\frac{1}{\beta}\bar{r}^{(i)}-\bar{\text{KL}}^{(i)}$:
$\begin{split}\nabla\mathcal{J}_{\text{TB}}(\mathbf{B};\theta)=\frac{1}{BK}\sum_{i=1,j=1}^{i=B,j=K}A^{(i,j)}\nabla\log\pi_{\theta}(\mathbf{y}^{(i,j)}\mid\mathbf{x}^{(i)}),\end{split}$ (12)
where $A^{(i,j)}=(r^{(i,j)}-\bar{r}^{(i)})-\beta(\hat{\text{KL}}^{(i,j)}-\bar{\text{KL}}^{(i)})$. In sum, TBA appears to update model weights using a REINFORCE algorithm with a mean-reward baseline [williams1992simple] and a KL-divergence-regularized reward [ziegler2019fine]. However, note that $\hat{\text{KL}}^{(i,j)}$ is only an estimate of the KL divergence in the on-policy case because off-policy data is not sampled from $\pi_{\theta}$ by definition, causing loss of equivalence to mean-baseline REINFORCE with a KL-regularized reward in the asynchronous settings that TBA operates in. Deviating further, TBA may also reset the reference policy similar to liu2025prorl.
Appendix B TBA Experiment Details
@@
For MR and PFT, we implement TBA by building on the RLOO trainer class of the Hugging Face TRL library [vonwerra2022trl]. For MR and PFT, we make modifications to the baseline hyperparameters as shown in Table 2 and Table 3, respectively.
Our RT implementation of TBA built on the TB trainer used in lee2024learning, and we largely utilize their code’s default hyperparameters, with differences noted in Appendix B.3.
B.1 GSM8K Mathematical Reasoning (MR)
All baselines and TBA results use the same starting point, a RhoMath-1B [lin2024rho] model SFTed on GSM8K training data by kazemnejad2024vineppounlockingrlpotential – “realtreetune/rho-1b-sft-GSM8K” on Hugging Face. The baseline model achieves 40.340.3% test accuracy. For the VinePPO baseline, training time is estimated using their reported 380 seconds per training step and 650 steps for GSM8K training [kazemnejad2024vineppounlockingrlpotential]. We implement the GRPO baseline using its Hugging Face TRL trainer, see Appendix E.1 for details.
Table 2: TBA Training Hyperparameters for the GSM8K MR task.
@@ B.1 GSM8K Mathematical Reasoning (MR)
Most-On-Policy Sampling Probability 0.95 mm in Section 5.4
@@
In Figure 1, TBA uses the settings in Table 2 with a few modifications to shorten training time by an additional 3030%, down to 82 minutes on 4xA100 GPUs. In particular, we observed in initial testing (see Appendix E) that using the hyperparameters in Table 2 led to no improvement in performance for the final 300 steps. Thus, we shrank the training duration from 1000 steps to 700 (98000 episodes with batch size 140). Additionally, we made the following modifications to attempt to reduce the variance of the shortened run: 350 stable learning rate steps (down from 450), 0.014 Beta Initial Value (up from 0.012). We ran this experiment three times – obtaining performances of 55.855.8%, 53.953.9%, and 54.154.1% – and reported the mean accuracy 54.654.6% in Figure 1.
@@
The 700-step result we show in Figure 1 has standard error 0.60.6%, which is a little more variance than what we observe in the original 1000 step setup (the blue line in the bottom left plot of Figure 12 shows the mean and standard errors for the 1000 step runs). Future work could further explore variance reduction approaches/hyperparameters for shorter runs (and for TBA/RL more generally).
One way to deal with variance is to choose a checkpoint based on a held out validation set, a strategy used to produce the VinePPO result [kazemnejad2024vineppounlockingrlpotential]. We do not use this approach but note that our results would likely benefit significantly ($\approx 1$%) from it. In particular, each of our runs tends to produce (at some point during training) a higher performance than the performance at the final step – this is expected if you consider that the model performance varies around the average value it converges to towards the end of training (e.g., see again the blue line in the bottom left plot of Figure 12). Despite its being lower than the maximum performance achieved at any point during training, we report this final step performance, which is what a user of TBA could expect to obtain without applying techniques like early-stopping.
@@
Most-On-Policy Sampling Probability 0.5 mm in Section 5.4
@@
B.2 TL;DR Preference Fine Tuning (PFT)
@@
All of our PFT results were run in 32-bit precision and without DeepSpeed, which was used by baselines we compared against [noukhovitch2024asynchronous]. Lower precision and such training acceleration packages could further improve the speedups we show. Relatedly, for our 2.82.8B runs, we used gradient checkpointing to fit more examples into a micro batch, which led to slowdowns at this scale (i.e., we only have a 3.83.8x speedup over the baseline in this setting). We leave the optimization of our framework with appropriate packages to future work. Finally, we used a large number (10000) of initial completions in the buffer, and future work should confirm that a smaller number (e.g. 1000) works equally well – note that a small number worked well for GSM8K.
B.3 Automated Red Teaming (RT)
Unlike our MR and PFT implementations, our RT implementation uses the code of lee2024learning as a baseline and thus does not follow the Hugging Face TRL trainer style. We discuss hyperparameters in the context of their trajectory balance trainer approach below. We adopt their code largely as it is, with the exception that our TBA implementation uses larger replay buffers than lee2024learning to accommodate our scaled search for trajectories. Additionally, unlike lee2024learning, we remove the oldest samples when the buffers fill up as opposed to the lowest reward samples.
For the Llama results in Table 6, our hyperparameter choices largely follow those of lee2024learning. We train for 5000 steps with batch size 128. For temperature sampling, we use a low and high of 0.7 and 2.0, respectively. We use a reward schedule horizon of 1000. The language model schedule end is 1.2, and its schedule’s horizon is 2000. We prioritize sampling based on reward. We use Beta 0.05. We use sync period ($k$) 10. We use 6 searchers. We use most-on-policy probability ($m$) 0.5 and 0.6 in combination with maximum buffer sizes 150000 and 130000, respectively. By having a smaller maximum buffer size and larger $m$, the latter setting is expected to focus on more recently generated data and prioritize reward over diversity, which is what we observe in Table 6.
For the GPT2 results in Figure 5 and Table 6, our hyperparameter choices again largely follow those of lee2024learning. We train for 5000 steps with batch size 128. We use Beta 0.05. We use sync period ($k$) 10. We use most-on-policy probability ($m$) 0.5. We cap the maximum buffer size at 100000 in all experiments; we additionally prevent the buffer from growing past its size as of step 4000 to encourage focusing on more recent data (larger most on policy probabilities $m$ may provide a similar effect). We test searcher counts 2, 6, 14, 30, 62, and we use 3 runs per configuration for error bars and averages.
@@ B.3 Automated Red Teaming (RT)
In Table 6, there is a slowdown as the number of searchers scales. We believe this is largely addressed by a newer version of our code that uses more efficient buffer communication, but we have not re-run these results yet to confirm this. In any case, developing more efficient TBA implementations is an interesting direction for future work given TBA’s ability to effectively leverage large-scale data generation.
Appendix C TBA Experiment Details
Our TBA experiments train Qwen 2.5 7B Base [qwen25] on MATH [hendrycksmath2021] or Qwen 3 4B Instruct 2507 [yang2025qwen3] on 9000 generated Countdown problems [stojanovski2025reasoninggymreasoningenvironments], testing on MATH-500 or 1000 held-out Countdown problems, respectively.
In Figure 6, we compare TBA to the Dr. GRPO [liu2025understanding] implementation in PRIME-RL [primeintellect2025primeRL]. The hyperparameters are the same for both methods and given in Table 4. Note that TBA is the same as Dr. GRPO when $\rho=1$ (and there’s no asynchrony) or when $\beta=0$. For more details on TBA and ablation studies, please see Appendix D.
Table 4: Training Hyperparameters for the MATH-500 and Countdown tasks. For Countdown, we train for up to 500 steps instead of 400. Also, Countdown experiments use Qwen 3 4B Instruct 2507 instead of Qwen 2.5 7B Base.
Hyperparameter Value Reference
Model Qwen 2.5 7B Base
Learning Rate $1\times 10^{-6}$
Generation Temperature 1.0
Max Prompt Sequence Length 1024
Max Model Length 3072
Number of Prompts per Batch 32
Number of Completions per Prompt 16 $K$ in Section 4.1
Batch Size (effective) 512
Number of Training Steps 400
Total Prompts Seen 12800
Total Episodes 204800
CISPO IS lower/upper bound 0/8
TBA-specific hyperparameters
Beta (KL coefficient) 0.005 $\beta$ in Equation 5
Reference policy reset period 50 $\rho$ in Section 6
Appendix D TBA Ablations and Contextualization in RL Landscape
TBA adds importance sampling (IS) to TBA with CISPO-style clipping of the IS ratio [chen2025minimax]. As shown in Figure 7, some tasks benefit from IS, though it is not critical for all tasks. TBA also resets the reference policy instead of reducing $\beta$ throughout training. Figure 8 shows that resetting the reference policy is critical for obtaining high performance, having an effect on accuracy similar to IS’s.
Figure 7: We add clipped importance sampling to TBA due to its helpfulness on some tasks.
Figure 8: We show the effect of turning off resetting and the effect of turning off clipped importance sampling in TBA.
Notably, resetting the reference policy too frequently risks removing the benefits of KL regularization. This reset period $\rho$ is thus a key hyperparameter for TBA (along with $\beta$): larger values of both $\rho$ and $\beta$ are expected to increase the optimization pressure toward the original reference policy, while smaller values allow freer optimization of the task reward but potentially risk stability. We emphasize that our Dr. GRPO baseline is exactly TBA when $\beta=0$, and this baseline displays lower robustness to off-policy data (see Figure 6).
Figure 9: When replacing importance sampling clipping with IcePop masking, TBA is still effective.
Instead of clipping the IS ratio, recent studies have shown that masking the gradient for tokens with large or small ratios can be effective [IcePop2025, zheng2025prosperity]. In Figure 9, we show that the TBA approach to KL regularization (resetting every 50 steps, adding the log probability ratio to the reward) is effective when used with such masking (specifically IcePop’s) instead of with CISPO-style clipping. Critically, such masking is not sufficient to gain robustness to highly off-policy data when used without TBA’s approach to KL regularization, as shown in Figure 10.
Figure 10: IcePop masking alone is not sufficient to reach the off-policy robustness of TBA.
The most similar approach to TBA’s that we are aware of is that of Kimi K1.5 and K2 [team2025kimi, team2025kimi2]. As shown in team2025kimi2, the Kimi K2 log probability ratio utilizes probabilities from the training policy and the inference policy: this is similar to KL regularization that resets the reference policy every step to be the policy used at inference time. Notably, to the extent that staying closer to the original reference policy is beneficial, this approach may underperform due to its frequent resets. While the K2 approach has no ability to alter the reset period and use a reference policy different from the inference policy, TBA has the flexibility to use longer/shorter reset periods. In our experiments, TBA resets the reference policy to match the current training policy just once every $\rho=50$ steps, potentially providing stronger adherence to the base policy. The other main difference is that TBA includes the mean log probability ratio inside the control variate, based on the principled derivation given in Section 3.2. Figure 11 shows that the Kimi approach and TBA have similar off-policy performance. Future work could investigate performance difference in settings where adherence to the base policy may be important (e.g. settings that induce reward hacking like cundy2025preference).
Figure 11: The Kimi K2 approach to KL regularization is similar to that of TBA and provides similar off-policy performance. After a small hyperparameter study, we use $\tau=0.005$ for K2.
Gradient Comparisons
The gradients of objective variants we consider are as follows. Note we hold constant the clipping strategy across methods, testing off-policy performance effects of key method aspects shown in red. Relatedly, in all cases, we use PRIME-RL’s default GRPO-style normalization when summing token-level gradients.
[…]
where $\lambda^{(j)}_{t}=\frac{\pi_{\theta}(y_{t}|x,y^{(j)}_{i<t})}{\pi_{\text{gen}}(y_{t}|x,y^{(j)}_{i<t})}$, $\Lambda^{(j)}_{t}=\frac{\pi_{\theta}(y_{t}|x,y^{(j)}_{i<t})}{\pi_{\text{ref}}(y_{t}|x,y^{(j)}_{i<t})}$, $\log\lambda^{(j)}=\sum_{t=1}^{|y^{(j)}|}\log\lambda^{(j)}_{t}$, $\log\Lambda^{(j)}=\sum_{t=1}^{|y^{(j)}|}\log\Lambda^{(j)}_{t}$, $\overline{\log\Lambda}$ is the mean of $\log\Lambda^{(j)}$ across $K$ sequences, and sg is the stop-gradient function. Normalization omitted for brevity. Finally, note that TBA resets $\pi_{\text{ref}}$ to be $\pi_{\theta}$ every $\rho$ steps.
brantley2025accelerating also use a regression objective similar to TBA’s, but instead learn offline a value function that plays the role of the control variate.
Appendix E TBA GSM8K Ablation Studies